A couple bought some stock for $40 per share that pays an annual dividend of $0.80 per share. After 1 year the price of the stock was $44. Find the simple interest rate on the growth of their investment.
how much % ?
If $5000 is invested at 7% annual simple interest, how long does it take to be worth $8150?
How many years?
Find the future value of an annuity of $1100 paid at the end of each year for 15 years, if interest is earned at a rate of 6%, compounded annually. (Round your answer to the nearest cent.)
How much $ ?
Find the future value of an ordinary annuity of $500 paid quarterly for 3 years, if the interest rate is 14%, compounded quarterly. (Round your answer to the nearest cent.)
How much $ ?

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

They started with $40 and at the end of 1 year had $0.80 of dividend and the stock was worth $44.
That is an increase of $4.80/$40 starting.
that is 12% increase.
======================
8150=5000(1+.07)^n
divide by 5000
1.63=(1.07)^n
ln both sides
0.48858=n ln (1.07)
divide ln (1.63) by ln (1.07) and round then.
7.22 years.
======================
FV=(1+r)*P(1+r)^n-1/r
=(1+(.06))*1100(1+.06))^15-1/0.06
=1.06*1.3965, round only at end.
=$27139.78
======================
FV=500(1+r/n))^4n-1/(r/n)
=500(1+.14/4)^12-1/(0.14/4)
=500(1.035)^12-1/0.035
=$7300.98

Answer by MathTherapy(10551)


A couple bought some stock for $40 per share that pays an annual dividend of $0.80 per share. After 1 year the price of the stock was $44. Find the simple interest rate on the growth of their investment.
how much % ?
If $5000 is invested at 7% annual simple interest, how long does it take to be worth $8150?
How many years?
Find the future value of an annuity of $1100 paid at the end of each year for 15 years, if interest is earned at a rate of 6%, compounded annually. (Round your answer to the nearest cent.)
How much $ ?
Find the future value of an ordinary annuity of $500 paid quarterly for 3 years, if the interest rate is 14%, compounded quarterly. (Round your answer to the nearest cent.)
How much $ ?
I'll do 1 for you. You have to do MOST of the work, or if I do all, what are you going to do?
2) It'll take 9 years for $5,000 to accumulate to $8,150.
However, if the money is left in an account and is not touched, the interest will be COMPOUNDED, and therefore, it'll take less time to accumulate to $8,150.